OpenStudy (anonymous):

Use the problem below to answer the question: When a current is passed through a solution of salt water, sodium chloride decomposes according to the following reaction: 2NaCl + 2H2O → 2NaOH + Cl2 + H2 How many grams of chlorine gas, Cl2, are given off if 7.5 grams of sodium chloride, NaCl, are decomposed? The molecular mass of NaCl is 58.44 g/mole. The molecular mass of Cl2 is 70.90 g/mole. Using these values, finish the above problem. _______ grams of chlorine gas

OpenStudy (anonymous):

@apoorvk sorry to bother you but can you help me with solving this equation

OpenStudy (chmvijay):

2*58.44 gram of NaCl gives =70.9 gram of chlorine 7.5 gram gives=(70.9*7.5) /2*58.44 =4.54 gram of chlorine

OpenStudy (anonymous):

thanks but can you like run this whole problem down to me...I have a quiz soon